### v1

# Broadband Two-Dimensional Far-Field Beam Profiling of Commercial CW Terahertz Photomixers from 0.2 to 1.5 THz

Abstract
We present a systematic far-field characterization of four commercial PIN diode terahertz photomixers over the 0.2–1.5 THz frequency range using a two-dimensional (2D) raster scanning method. The emission pattern of each transmitter, equipped with an integrated hyper-hemispherical silicon lens, was characterized using a broadband Schottky diode receiver over a 35$\times$35 mm2 grid. The results reveal consistent frequency-dependent beam divergence and the emergence of distinct Airy diffraction patterns at intermediate frequencies, attributed to lens-induced aperture effects. This comprehensive mapping underscores the importance of full 2D beam profiling for understanding THz propagation and offer insights into lens design and photomixer packaging for optimized system performance.
keywords:
Terahertz , Photomixer , Frequency-Domain Spectroscopy , Beam Profiling
journal: Infrared Physics & Technology
[…]
[1]organization=Department of Materials and Production, Aalborg University, city=Aalborg East, postcode=DK-9220, country=Denmark
1 Introduction
Photomixers are a well-established technology for generating continuous-wave (CW) terahertz (THz) radiation, offering broad frequency tunability and room-temperature operation. These attributes make them highly attractive for applications in imaging, spectroscopy, and high-speed wireless communications. However, a critical yet often underexplored aspect of photomixer performance is the spatial emission pattern of the THz beam, which directly affects coupling efficiency, optical alignment, and the integration of photonic and quasi-optical components in THz systems.
Recent work in CW-THz imaging has highlighted the importance of detailed beam characterization, particularly using coherent detection techniques capable of recovering both amplitude and phase [1]. These methods, based on frequency scanning and fringe analysis, have been successfully used to reconstruct complex THz fields, including those shaped by integrated silicon lenses. Despite their strengths, such approaches require phase-stable interferometric setups, are sensitive to alignment, and are limited by very slow acquisition speeds.
Only a few other studies have explored the far-field behavior of PIN diode THz photomixers in commercial THz frequency-domain spectroscopy (THz-FDS) systems [2, 3]. However, these studies relied on one-dimensional (1D) angular scans in the principal E- and H-planes [2, 3]. While informative, these 1D measurements capture only partial information and may not disclose key features of the full beam profile, including asymmetries, off-axis lobes, and diffraction-induced ring structures. These effects become especially pronounced when silicon lenses are integrated with the transmitter, as they introduce complex, frequency-dependent propagation behavior. In the far-field regime—defined as the region where the electromagnetic field becomes predominantly transverse and angular distribution becomes independent of distance—the spatial beam characteristics can be meaningfully interpreted and compared. This regime begins when:
$r\geq\frac{2D^{2}}{\lambda}=\frac{2D^{2}\nu}{c}$ (1)
where $r$ is the distance between the transmitter and receiver plane, $D>\lambda/2$ is the effective aperture diameter and $\lambda$ is the free-space wavelength [4]. For photomixers with integrated silicon lenses, the high refractive index of silicon ($n_{\mathrm{Si}}=3.42$) shortens the wavelength inside the material and extends the near-field region [3], complicating the interpretation of beam patterns and potentially contributing to the broadening and structured features observed in prior measurements. Consequently, at frequencies above 0.3 THz, the near-field region can extend beyond the silicon lens, which Smith et al. [3] suggested may explain the beam broadening and additional features observed in their measurements.
In this study, we address these limitations by presenting comprehensive two-dimensional (2D) far-field intensity maps of four nominally identical photomixers across a broad frequency range (0.2–1.5 THz). Using an incoherent single-pixel detection scheme based on lateral scanning of a broadband Schottky diode receiver, we directly measure the intensity distribution of the emitted THz beam without requiring phase sensitivity. This approach enables robust and high-resolution mapping of key spatial features—including beam divergence, asymmetries, and frequency-dependent diffraction effects—that are often inaccessible in 1D scans. Our results reveal consistent patterns across devices and frequencies, including clear Airy-like structures resulting from aperture-limited propagation through the silicon lens. These measurements provide new insight into the optical behavior of photomixer-based THz sources and offer valuable benchmarks for transmitter modeling, alignment strategies, and quasi-optical system design.
2 Experimental Work
The measurements were performed with a TeraScan 1550 system (Toptica Photonics) in combination with an incoherent Schottky diode receiver (Toptica TD-RX-1), which offers a bandwidth of 50 GHz–1.5 THz and a noise-equivalent power of 70 pW/$\sqrt{\mathrm{Hz}}$ at 100 GHz and 1000 pW/$\sqrt{\mathrm{Hz}}$ at 1 THz. The receiver was mounted on a motorized XY-stage to perform raster scanning with a 1 mm step size, enabling full two-dimensional mapping of the far-field emission from THz photomixers. The photomixers themselves were mounted on a Z-stage, allowing the distance to the image plane to be adjusted between 90 and 150 mm. The experimental setup is illustrated in Fig. 1(a).
Figure 1: (a) Illustration of the experimental setup and the bowtie antenna design of the THz photomixer transmitters. The abbreviations are: tx - photomixer transmitter; rx - Schottky diode receiver; WG - waveguide; PD - PIN photodiode. The antenna design inset is adapted from Ref. [2], © 2021 Optical Society of America. (b) Power spectra of the four THz photomixer transmitters, adapted from manufacturer data sheets.
To ensure far-field conditions across all frequencies up to 1.5 THz, the transmitter-to-receiver distance in our setup exceeds 35 mm, satisfying the far-field criterion even with the presence of an integrated silicon lens. Previous studies [2, 3] have predominantly utilized angular scanning to characterize one-dimensional beam profiles in the E- and H-planes. These measurements were performed by rotating the transmitter relative to the receiver while maintaining polarization alignment. Angular scans are particularly advantageous for divergent beams, as they preserve a constant transmitter-receiver distance and angular accuracy. However, they inherently yield only 1D cross-sections of the beam. Although combining multiple angular scans can reconstruct a 2D radiation pattern, this approach increases experimental complexity and complicates data interpretation. Moreover, even single-plane profiles may produce ambiguous results due to the lack of comprehensive spatial information and overlook beam asymmetries or off-axis features. In contrast, our study employs a full 2D lateral scan in the far field, enabling a more complete spatial characterization of the THz beam. While slight distortions are expected due to small variations in propagation distance and the angular sensitivity of the receiver, the resulting 2D profiles provide a significantly more informative picture. For instance, when the transmitter is positioned at $z=$ 100 mm from a 50 mm $\times$ 50 mm image plane, the optical path length to the outermost pixels deviates by only a few percent relative to the center. For each measurement, the transmitter was aligned such that the THz electric field was vertically polarized (E-plane parallel to the $x$-axis) and positioned in front of the Schottky diode at distance $z$.
We characterized four nominally identical InGaAs PIN diode THz photomixers featuring an extended bowtie antenna geometry (See inset in Fig. 1(a)) with an effective aperture diameter of approximately $D\approx 1$ mm. The power spectra of these devices, adapted from test reports provided by the Fraunhofer Heinrich Hertz Institute (HHI), are shown in Fig. 1(b). All four antennas perform within specified parameters; however, one device (indicated by the blue line) clearly outperforms the others in terms of output power. This particular unit (TC190201) originates from a newer production batch. Despite this, all four photomixers are based on the same chip design, and according to the manufacturer, no changes have been made to the fabrication or assembly processes.  [^1] According to HHI, based on private correspondence with Dr. Anselm Deninger, Director Technical Sales Support, Toptica Photonics AG.
3 Results and Discussion
First, we evaluated the measurement repeatability of our setup. Ten successive measurements were performed on a single antenna at $z=145$ mm, and the resulting means and standard deviations for the E- and H-planes are shown in Fig. 2(a). Overall, the repeatability is good, with the largest variance occurring near the main peak. Moreover, the general beam profiles in both the E- and H-planes are in good agreement with previously published results [3, 2].
Figure 2: Measurement repeatability test at $\nu=0.5$ THz using the TC190201 antenna (a), and evaluation of the influence of standing waves in the setup (b) and (c).
Standing waves between the transmitter and receiver can affect measurements when using CW THz radiation, and must therefore be considered. To assess their influence, we scanned the receiver position over a 300 $\mu$m interval in steps of 50 $\mu$m between consecutive beam profile measurements at 0.5 THz. This interval corresponds to approximately $\lambda/2$. The results for the E- and H-planes are shown in Fig. 2(b) and (c), respectively. A noticeable decrease in the peak amplitude is observed at $\Delta z=200$ and 250 $\mu$m, which may indicate the presence of standing wave effects when the receiver is positioned directly in front of the transmitter.
The radiation patterns were measured over a 35 mm $\times$ 35 mm area in 1 mm increments for frequencies ranging from 0.2 THz to 1.5 THz, with the receiver positioned at $z=90$ mm in front of the antennas. The resulting beam profiles for the four transmitters are presented in Fig. 3. For consistency, the color scale in each map is normalized to the output of the TC190201 antenna at the corresponding frequency.
Figure 3: Radiation patterns of the (a) TC190201, (b) TC190106, (c) TC180501, and (d) TC170806 antennas at frequencies from 0.2 to 1.5 THz. Each map covers a 35 mm $\times$ 35 mm area in 1 mm increments and is measured at a distance $z=90$ mm in front of the antennas.
At a first glance, the beam diameter decreases from 0.2 THz, reaching a minimum around 0.6$\sim$0.9 THz, before increasing again up to 1.5 THz. Additionally, the radiation patterns exhibit irregularities particularly around 0.4$\sim$0.5 THz and 1$\sim$1.2 THz. At higher frequencies, the patterns appear increasingly blurred, which can be attributed to a reduced signal-to-noise ratio and the fixed 1 mm spatial resolution. The observed asymmetries in the E- and H-plane profiles have previously been attributed by Nellen et al.[2] to the asymmetric feed point structure of the antenna, as shown in Fig. 1(a). Notably, the dimensions of the bracket structure correspond to $\lambda/4$ at 0.35 THz, and the extended bowtie geometry has been proposed to act as a broad dipole, generating frequency-dependent side lobes. A closer examination of the beam profiles in the 0.5–1 THz range reveals the presence of weaker rings surrounding the main lobe. In Fig. 4(a)-(d), the dynamic range at 0.7 THz is calculated for each antenna to highlight the radiation pattern.
Figure 4: Dynamic range of the emission patterns (35 mm×$\times$35 mm in 1 mm increments) measured at $z=90$ mm: (a)-(d) Of all four transmitters as measured at $\nu=$0.7 THz. (e)-(h) Of the TC190201 antenna at frequencies $\nu=$ 0.2, 0.4, 0.6 and 0.8 THz. And (i) estimated Airy disk radius as a function of frequency of the four THz transmitters. The solid black curve is the theoretical Airy disk radius assuming a 10-mm aperture radius. The dotted vertical line represent the maximum frequency, for which at least five measurements across the Airy disk can be obtained using 1 mm increments.
In all cases, an Airy-like pattern is observed—characterized by a bright central spot (Airy disk) surrounded by concentric bright rings—arising from diffraction through a circular aperture. Notably, the Airy disk appears approximately the same size across all antennas. Figure 4(e)-(h) shows the dynamic range maps of the TC190201 antenna across the 0.2–0.8 THz frequency range. Here, the diameters of both the central spot and the concentric rings are seen to decrease with increasing frequency. This behavior may be attributed to diffraction of the THz waves by the circular aperture of the integrated silicon lens. The minima of an Airy pattern observed far from the aperture occur at angles given by
$\sin\theta\approx m\frac{\lambda}{d},$ (2)
[…]
To estimate the Airy disk radius from the measured beam intensity profiles, we fit the normalized intensity data along a single spatial dimension (e.g., $x$, $y$, diagonal, and anti-diagonal) to a theoretical Airy pattern given by
$I(r)=\left[\frac{2J_{1}\left(\alpha\,r\right)}{\alpha\,r}\right]^{2},$ (3)
where $r$ is the radial coordinate, $J_{1}$ is the first-order Bessel function of the first kind, $\alpha=\pi d\nu/cz$ with $c$ the speed of light, and $z$ the propagation distance from the circular aperture to the detection plane [5]. In practice, we fit the experimental data to this function using $\alpha$ as a free parameter. The first minimum of the Airy pattern, which defines the Airy disk radius $R$, occurs where $J_{1}(\alpha\,r)$ first vanishes. Since the first zero of $J_{1}(x)$ occurs at $x\approx 3.8317$, we obtain the relation $R=3.8317/\alpha$. This constant thus directly links the fitted parameter $\alpha$ to the physical Airy disk radius, providing a reliable means to extract the diffraction-limited spot sizes from the experimental data. To improve the robustness of the estimate, we perform the fit along the $x$, $y$, diagonal, and anti-diagonal directions, and then calculate the mean value of the resulting fits. The results for the four THz transmitters between 0.2 and 0.8 THz are shown in Fig. 4(i). The solid black curve represents the theoretical Airy disk radius
$R(\nu)=\frac{3.8317\,c\,z}{\pi d\nu},$ (4)
[…]
4 Conclusion
We characterized four nominally identical, commercially available PIN diode THz photomixers featuring an extended bowtie antenna structure. This was accomplished by mapping their two-dimensional far-field radiation patterns across a range of frequencies. Despite noticeable variations in the radiation patterns among individual devices, a consistent overall trend emerged: A dominant central spot surrounded by weak concentric rings. Our analysis indicates that these Airy-like patterns most likely result from diffraction of the THz waves by the circular aperture defined by the silicon lenses. In principle, such diffraction caused by the silicon lens could be mitigated by etching micro-structured lenses directly onto the backside of the antenna substrate [6, 7].
Acknowledgment
We gratefully acknowledge Dr. Anselm Deninger, Director of Technical Sales Support at TOPTICA Photonics AG, for his insightful discussions, for providing valuable details on the photomixer antenna design, and for sharing the calibrated power spectra of the photomixers.
Funding
This work was supported by the Innovation Fund Denmark Grand Solutions program (grant no. IFD-7076-00017B).
Disclosures
The authors declare no conflicts of interest.